Net Present Value—Unequal Lives

Project 1 requires an original investment of $94,300. The project will yield cash flows of $17,000 per year for eight years. Project 2 has a calculated net present value of $23,200 over a six-year life. Project 1 could be sold at the end of six years for a price of $75,000.

Use the Present Value of $1 at Compound Interest and the Present Value of an Annuity of $1 at Compound Interest tables shown below.

Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162
Present Value of an Annuity of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 1.833 1.736 1.690 1.626 1.528
3 2.673 2.487 2.402 2.283 2.106
4 3.465 3.170 3.037 2.855 2.589
5 4.212 3.791 3.605 3.352 2.991
6 4.917 4.355 4.111 3.784 3.326
7 5.582 4.868 4.564 4.160 3.605
8 6.210 5.335 4.968 4.487 3.837
9 6.802 5.759 5.328 4.772 4.031
10 7.360 6.145 5.650 5.019 4.192

a. Determine the net present value of Project 1 over a six-year life with residual value, assuming a minimum rate of return of 10%. If required, round to the nearest dollar.
$

b. Which project provides the greatest net present value?
Answer #1
Calculation of Net present value
Present value of cash inflow = cash inflow*PVIF @10%,6yr
present value of cash inflow = 17000*4.355
present value of cash inflow = 74035
Present value of salvage value = 75000* 0.564 (i.e 10%,pv ,6yr)
Present value of salvage value = 42300
Net present value = present value of cash inflow-initial investment
Net present value = (74035+42300)-94300
Net present value = $22035
b) Project 2 Provide greatest Net present value